OpenStudy (anonymous):

In a heptagon QRSTUVW, angle Q measures 130*, angle R measures 140*, angle S and angle T measures 160* each, angle U measures 100* & angle V measures 90*, what is the measure of angle W?

OpenStudy (anonymous):

Sum of internal angles of a heptagon is 900 degrees. So, W = 900 - (130+140+160+160+100+90) = 900 - 780 = 120 degrees 2) A polygon with all sides the same length and all angles of equal measure

OpenStudy (anonymous):

Thank you ;D

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

:)